A baseball is thrown horizontally with an initial velocity of 20 m/s from


the height of 1.5 m. How long will it take to reach the ground? How far


will it go horizontally?

1
Expert's answer
2022-03-03T12:05:03-0500

Time in air

"t=\sqrt{2h/g}=\sqrt{2*1.5/9.8}=0.55\:\rm s"

The horizontal distance

"l=v_0t=20*0.55=11\:\rm m"